Use the formula S = √(30Df ) to determine the distance required for a car to skid to a stop. (S is the car's in miles per hour, D is the length of the skid marks in feet, and f is the coefficient of friction between the tires and the road.)
f = 0.8 ; S = 45 mi/hr
NOTE:
√(30Df ) is a square root of 30Df

Answer by rothauserc(4718)

S = sqrt(30DF)
:
square both sides of =
:
S^2 = 30DF
:
D = S^2 / 30F
:
D = 45^2 / (30*0.8) = 84.375 feet